Directions: Sp;ve the equation by completing the square. Express any complex numbers using i notation.
x^2+8x=3

Answer by Nate(3500)

x^2 + 8x = 3
(x + 4)^2 = 19
x + 4 = +- sqrt(19)
x = -4 +- sqrt(19)